# Evidence for Low Universal Equilibrium Black Hole Spin in Luminous Magnetically Arrested Disks

### v1

@@ Abstract
Relativistic collimated outflows, or jets, provide a crucial mode of active galactic nucleus feedback. Although the jets extract their energy from the black hole (BH) rotation, their effect on the BH spin is poorly understood. Because the spin controls radiative and mechanical BH feedback, lack of first-principles models for BH spin evolution limits our ability to interpret observations, including the recent LIGO-Virgo-KAGRA spin constraints. Particularly important are luminous disks, which rapidly grow and strongly torque their BHs. Jet-less and weakly magnetized “standard” luminous disks spin-up their BHs to near-maximum dimensionless spin, $a_{\text{eq,NT}}=0.998$. However, sufficient large-scale vertical magnetic flux can cause the inner disk to enter a magnetically arrested disk (MAD) state, whose jets can efficiently extract BH rotational energy and significantly spin-down the BH. Indeed, Lowell et al. (2024) found that non-radiative MADs spin-down their BHs to very low $a_{\text{eq,MAD}}^{\text{nr}}=0.07$. Moreover, their analytic model predicted that luminous MADs also spin-down their BHs to low $a_{\text{eq,MAD}}^{\text{lum}}\sim 0.3-0.5$. To test this prediction, we perform 3D general relativistic (radiation) magnetohydrodynamic (GR(R)MHD) simulations of MADs across a wide range of BH spin ($-0.9\leq{}a\leq 0.99$) and disk dimensionless thickness ($0.03\leq{}h/r\leq 0.3$, which corresponds to Eddington ratio, $0.35\leq{}\dot{m}/\dot{m}_{\text{Edd}}\leq\infty$). We find that luminous MADs spin-down their BHs to a low universal equilibrium spin value, $a_{\text{eq,MAD}}^{\text{lum}}\approx 0.3$ for $0.03\leq{}h/r\leq 0.1$. Moreover, we find evidence for quadratic convergence, $a_{\text{eq,MAD}}\approx 0.31-2.71(h/r)^{2}\to 0.31$ as $h/r\to 0$. We attribute this to disk thermodynamics becoming irrelevant as the cooling becomes more aggressive and magnetic forces start to dominate. We finish by discussing the astrophysical implications.
Black holes — relativistic jets — black hole spin – magnetohydrodynamics
1 Introduction
@@
Relativistic jets are observed across a range of astrophysical objects, including active galactic nuclei (AGN), ranging from low luminosity AGN (e.g., M87) to luminous quasars (Kong & Ho, 2018), X-ray binaries (XRBs, Done et al., 2007), tidal disruption events (TDEs, Burrows et al., 2011; Bloom et al., 2011; Tchekhovskoy et al., 2014), and gamma ray bursts (GRBs) powered by collapsars (Woosley, 1993; Goldstein et al., 2016) and neutron star (NS) mergers (Foucart et al., 2015; Metzger, 2019; Nakar, 2020).
@@
A popular mechanism for jet launching is the Blandford-Znajek (BZ) mechanism (Blandford & Znajek, 1977). In this mechanism, large-scale magnetic flux taps into the spin energy of the black hole (BH) and leads to BH spin-down. The jet power scales approximately as $P_{\rm jet}\sim\Phi^{2}_{\rm BH}a^{2}$, where $\Phi_{\rm BH}$ is the magnetic flux on the BH, and $-1\leq a\leq 1$ is the dimensionless BH spin (Blandford & Znajek, 1977; Komissarov, 1999; Tchekhovskoy et al., 2010).
@@
Since BH spin controls jet power, it is crucial to understand what determines the spin across different astrophysical sources. However, observationally constraining the BH spin is challenging and often subject to tension, as different measurement methods yield inconsistent spin distributions. Reflection methods in XRBs and AGN consistently measure moderate to high BH spins, $a\geq 0.5$ (Reynolds, 2021; Draghis et al., 2024). In contrast, gravitational wave observatories such as LIGO-Virgo-KAGRA (LVK) consistently derive small spins, $a\leq 0.4$, from waveform data (Abbott et al., 2019; Wysocki et al., 2019; LIGO Scientific Collaboration et al., 2023; Edelman et al., 2023). Constraining and understanding spin evolution is crucial for revealing the formation pathways of LVK sources and advancing our knowledge of binary evolution, because as much as $\sim 50\%$ of LVK sources can come from stellar binary systems. (Edelman et al., 2023).
Due to the no-hair theorem, BHs cannot hold onto the magnetic flux by themselves. Matter in the form of an accretion disk needs to force the magnetic flux into the BH and keep it there. Accretion flows can directly affect the BH spin: the accreted material adds angular momentum to the BH and leads to its spin-up. Indeed, in the Novikov-Thorne (NT) accretion disk model (Novikov & Thorne, 1973), which does not include large-scale magnetic fields, the BH spins up due to accreting (Keplerian) disk’s angular momentum (Bardeen, 1970; Moderski & Sikora, 1996).
@@
In jetted accretion systems, where a large scale vertical magnetic field is present, spin evolution will depend on the interplay of accretion and ejection processes near the BH event horizon. Tchekhovskoy et al. (2010) found that the jet shape determines the jet angular momentum and power fluxes and thus the spin evolution of the BH (see also Blandford & Znajek (1977)). Since the disk dimensionless thickness, $h/r$, or the ratio of disk height to radius, is connected to the jet opening angle, the diet of the BH affects spin evolution. Thus, the BH spin extraction can depend on the disk geometric properties, which are determined by the disk cooling regime.
@@
The accretion rate, $\dot{m}$, and how close it is to Eddington value, $\dot{m}_{\rm Edd}$, at which the radiation becomes as important as gravity, roughly sets the disk cooling regime (e.g., Yuan & Narayan, 2014). In the highly super-Eddington regime ($\dot{m}\gg\dot{m}_{\rm Edd}$), the disk is radiatively-inefficient and photons are trapped in the flow and advected in with the accretion flow or out with the outflow (Popham et al., 1999; Wyithe & Loeb, 2012; Sądowski & Narayan, 2015a; Inayoshi et al., 2016). This results in a hot accretion flow, which takes the form of a geometrically-thick accretion disk. In numerical simulations, such disks typically have a thermal scale height of $h/r\simeq 0.3-0.5$. The extremely sub-Eddington regime, $\dot{m}\ll\dot{0}.01m_{\rm Edd}$, also creates a thick radiatively-inefficient accretion flow (RIAF), which takes the form of a thick disk (Narayan & Yi 1994, 1995a, 1995b). In the middle, $0.01\dot{m}_{\rm Edd}\lesssim\dot{m}\lesssim\dot{m}_{\rm Edd}$, lies the moderately sub-Eddington regime with an accretion disk that radiates efficiently and cools to thinner scale height values, which can become very small, $h/r\ll 1$.
@@
Observations have suggested that the Eddington ratio of the BH correlates with the XRB spectral state (Esin et al., 1997, 1998; Qiao & Liu, 2013; Gilfanov, 2010). The so-called soft state is believed to indicate a razor-thin disk, where the inner edge of the disk is expected to approximately be at the innermost stable circular orbit (ISCO, Pringle & Rees, 1972; Shakura & Sunyaev, 1973a; Mitsuda et al., 1984; Marcel et al., 2019). The so-called low/hard state is linked to a thicker disk in the extremely sub-Eddington regime, with electrons that are much cooler than the ions (Marcel et al., 2018, 2019; Liska et al., 2022).
@@
AGN are also observed over a wide range of Eddington ratios. Generally, we expect the luminous disks to be thinner than the radiatively-inefficient disks, although the iron opacity bump and magnetic support may complicate the picture (e.g., Jiang & Blaes, 2020; Jiang & Dai, 2024; Hopkins et al., 2024; Squire et al., 2024; Kaaz et al., 2024). About $10\%$ of luminous quasars are radio-loud, which seems to be related to the strength of jets (Kellermann et al., 1989; Ivezić et al., 2002; Kellermann et al., 2016). Like XRBs, the measured AGN BH spins are often large. Black hole spin estimates from standard AGN SED continuum fitting methods assume a standard accretion disk and are sensitive to inclination angle (Hagen & Done 2023; see, e.g., Davis & Tchekhovskoy 2020, for limitations of the standard disk model in the context of AGN).
@@
More extreme accretion systems, such as collapsing massive stars producing BHs (collapsars) and binary compact object mergers, can achieve highly super-Eddington accretion rates and have thinner disks due to the effects of neutrino cooling. When neutrino luminosities are high, the disks can cool down to an aspect ratio of $h/r\sim 0.2$ (Chen & Beloborodov, 2007).
For most astrophysical disks sources, accretion can occur over a wide range of Eddington ratios, resulting in a variety of disk thermal thicknesses for the same type of astrophysical source, even varying throughout the evolution of the source in some cases. This variety of disk thickness values across different types of astrophysical sources motivates a systematic exploration of the effects that disk thickness has on BH spin evolution.
@@
For this, it is essential to capture the accretion and ejection physics, including the effects of strong magnetic fields, radiation, and curved spacetimes. This ambitious goal can be achieved with general relativistic (radiation) magnetohydrodynamic (GR(R)MHD) simulations. In this work, we focus on magnetically arrested disks (MADs, Igumenshchev et al., 2003; Tchekhovskoy et al., 2011), which are the natural end state of a disk with a sufficiently large vertical magnetic flux reservoir (Tchekhovskoy et al., 2011; Tchekhovskoy & Giannios, 2015; Jacquemin-Ide et al., 2021). Such highly magnetized disk states have also been found to emerge naturally due to dynamo action within the accretion disk (Liska et al., 2020; Jacquemin-Ide et al., 2024b).
@@
BH equilibrium spin is defined as the BH spin for which angular momentum lost through the jet balances the angular momentum and mass supplied by the disk. For thick MADs, GRMHD simulations revealed that the equilibrium spin is small, $a_{\rm eq}\simeq 0.035-0.07$ (Tchekhovskoy et al., 2012; Tchekhovskoy, 2015; Narayan et al., 2022; Lowell et al., 2024). Lowell et al. (2024) found that the BH only needs to accrete $20\%$ of its initial mass to spin down from $a=1$ to $a=0.2$. They also found that the BH loses most of its spin energy to the jets and that the positive angular momentum contribution from the accretion flow is suppressed by the large-scale Maxwell torques in the accretion disk (see also Manikantan et al. 2024). They then used this insight to derive a semi-analytic model for spin evolution in the thick nonradiative MAD regime. Using this model, they predicted that luminous, thin MADs would have low equilibrium spin values, $a_{\rm eq}\sim 0.3-0.5$. This is much lower than $a_{\rm eq}\sim 0.8$ obtained by Ricarte et al. (2023) based on their GRMHD simulations. Which is close to what is expected in NT disk ($a_{\rm eq}\sim 1$ and measured in weakly magnetized standard and normal evolution (SANE) accretion disk ($a_{\rm eq}\sim 0.9$) (Gammie et al., 2004; Chatterjee & Narayan, 2022).
@@
In this work we perform high-resolution 3D GRMHD simulations of cooled BH accretion disks in the MAD state to study BH spin evolution for thin accretion disks. This allows us to extend the work of Lowell et al. (2024) from the radiatively-inefficient thick to luminous MADs of varying disk scale height to physically interpret our simulation results.
@@
Section 2.1 presents the BH spin evolution equations. Section 2.3 describes our numerical setup and simulations. Section 2.4 explains how we calculate the torques on the BH. Section 3 presents our results, and, finally, Section 4 discusses the results and concludes. Throughout, we use units such that $G=c=1$. We adopt Lorentz-Heaviside units for the magnetic field, so that the magnetic pressure is given by $p_{\rm mag}=b^{2}/2$.
2 Methods and Equations
2.1 Spin evolution equations
Dimensionless BH spin is defined as
@@ 2.1 Spin evolution equations
where $J$ is the angular momentum and $M$ is the mass of the BH. The dimensionless spin spans $-1\leq a\leq 1$. The spin evolution of the BH can be computed via a set of ODEs for the specific energy and angular momentum of the BH (Moderski & Sikora, 1996; Lowell et al., 2024),
$dM=e_{\rm in}dm$ (2)
@@
$dJ=l_{\rm in}dm,$ (3)
where $l_{\rm in}$ and $e_{\rm in}$ are the specific angular momentum and energy fluxes, respectively, on the BH horizon, $r_{\rm h}=r_{\rm g}(1+\sqrt{1-a^{2}})$, and $dm$ is the amount of accreted mass. Here, $r_{g}=GM/c^{2}$ is BH gravitational radius. Combining Equations (2) and (3) gives the ODE for the change in BH spin with BH mass,
$\frac{da}{d\log M}=\frac{l_{\rm in}}{e_{\rm in}}-2a=\frac{s}{e_{\rm in}},$ (4)
where $\log$ is the natural logarithm, and we have defined the dimensionless spin-up parameter,
$s\equiv\frac{da}{dm}M=l_{\rm in}-2ae_{\rm in}.$ (5)
Computing the BH spin evolution is a question of determining, as a function of BH spin, the specific fluxes, $l_{\rm in}$ and $e_{\rm in}$, both of which depend on the (to be determined) accretion flow and jet structures.
2.2 Novikov-Thorne disk model
The NT disk model by Novikov & Thorne (1973) is an analytic model for a razor-thin accretion disk where the orbits are Keplerian, and the accretion mechanism is a viscosity of turbulent origin. From Bardeen (1970) and Moderski & Sikora (1996), the specific energy and angular momentum fluxes on the BH are,
$e_{\rm in}=\left(1-\frac{2r_{\rm g}}{3R_{\rm ms}}\right)^{1/2}$ (6)
and
$l_{\rm in}=\frac{2M}{3^{3/2}}\left(1+2(3R_{\rm ms}/r_{\rm g}-2)^{1/2}\right),$ (7)
where $R_{\rm ms}$ is the radius of the marginally-stable orbit, also known as the innermost stable circular orbit (ISCO). This radius is given by (Shapiro & Teukolsky, 1983),
$R_{\rm ms}/r_{\rm g}=3+Z_{2}-\left[(3-Z_{1})(3+Z_{1}+2Z_{2})\right]^{1/2},$ (8)

### v2

@@ Abstract
Relativistic collimated outflows, or jets, provide a crucial mode of active galactic nucleus feedback. Although the jets extract their energy from the black hole (BH) rotation, their effect on the BH spin is poorly understood. Because the spin controls radiative and mechanical BH feedback, lack of first-principles models for BH spin evolution limits our ability to interpret observations, including the recent LIGO-Virgo-KAGRA spin constraints. Particularly important are luminous disks, which rapidly grow and strongly torque their BHs. Jetless and weakly magnetized “standard” luminous disks spin up their BHs to near-maximum dimensionless spin, $a_{\mathrm{eq,NT}}=0.998$. However, sufficient large-scale vertical magnetic flux can cause the inner disk to enter a magnetically arrested disk (MAD) state, whose jets can efficiently extract BH rotational energy and significantly spin down the BH. Indeed, Lowell et al. [54] found that nonradiative, thick MADs spin down their BHs to very low $a_{\mathrm{eq,MAD}}^{\mathrm{thick}}=0.07$. Moreover, their analytic model predicted that luminous, thin MADs also spin down their BHs to low $a_{\mathrm{eq,MAD}}^{\mathrm{thin}}\sim 0.3\text{--}0.5$. To test this prediction, we perform 3D general relativistic (radiation) magnetohydrodynamic (GR(R)MHD) simulations of MADs across a wide range of BH spin ($0.9\leq{}a\leq 0.99$) and disk dimensionless thickness ($0.03\leq{}h/r\leq 0.3$, which corresponds to Eddington ratio, $0.35\leq{}\dot{m}/\dot{m}_{\mathrm{Edd}}\leq\infty$). We find that luminous, thin MADs ($0.03\leq{}h/r\leq 0.1$) efficiently spin down their BHs to a low universal equilibrium spin value, $a_{\mathrm{eq,MAD}}^{\mathrm{thin}}\approx 0.3$: a maximally spinning BH ($a=1$) spins down to $a=0.5$ after accreting just $25\%$ of its initial mass. Our results are consistent with quadratic convergence, $a_{\mathrm{eq,MAD}}^{\mathrm{fit}}\simeq 0.3-2.7(h/r)^{2}\to 0.3$ as $h/r\to 0$, which we attribute to the aggressive cooling that renders disk thermodynamics irrelevant and magnetic forces insensitive to thermal $h/r$. We finish by discussing the astrophysical implications.
I Introduction
@@
Relativistic jets are observed across a range of astrophysical objects, including active galactic nuclei (AGN), ranging from low luminosity AGN (e.g., M87) to luminous quasars [46], x-ray binaries [XRBs, 13], tidal disruption events [TDEs, 5, 4, 93], and gamma ray bursts (GRBs) powered by collapsars [98, 27] and neutron star (NS) mergers [22, 61, 64].
@@
A popular mechanism for jet launching is the Blandford-Znajek (BZ) process [3]: large-scale vertical magnetic flux taps into the spin energy of the black hole (BH) and leads to BH spin-down. The jet power scales approximately as $P_{\mathrm{jet}}\propto\Phi^{2}_{\mathrm{BH}}a^{2}$, where $\Phi_{\mathrm{BH}}$ is the vertical magnetic flux on the BH, and $1\leq a\leq 1$ is the dimensionless BH spin [3, 45, 94].
@@
Because BH spin controls jet power, it is crucial to understand what determines the spin across different astrophysical sources. However, observationally constraining the BH spin is challenging and often subject to tension, as different measurement methods yield incompatible spin distributions. Reflection methods in XRBs and AGN consistently measure moderate to high BH spins, $a\gtrsim 0.5$ [80, 14]. In contrast, gravitational wave observations by LIGO-Virgo-KAGRA (LVK) consistently prefer small spins, $a\lesssim 0.4$, from waveform data [48, 101, 49, 15]. Constraining and understanding spin evolution is crucial for revealing the formation pathways of LVK sources and advancing our knowledge of binary evolution, because up to $\sim 50\%$ of LVK sources can come from stellar binary systems [15]. Accretion flows can directly affect the BH spin: the accreted material adds angular momentum to the BH and leads to its spin-up. Indeed, in the Novikov-Thorne (NT) accretion disk model [73], which does not include large-scale magnetic fields, the BH spins up due to accreting (Keplerian) disk’s angular momentum [2, 63].
@@
Due to the no-hair theorem, BHs cannot hold onto the magnetic flux by themselves. Matter in the form of an accretion disk needs to force the magnetic flux into the BH and keep it there. Thus, in jetted accretion systems, where a large scale vertical magnetic field is present, spin evolution will depend on the interplay of accretion and ejection processes near the BH event horizon. Jet shape determines the jet angular momentum and power fluxes and thus the spin evolution of the BH [3, 94]. Since the disk dimensionless thickness, $h/r$, or the ratio of disk height to radius, is connected to the jet opening angle, the diet of the BH affects spin evolution. Thus, the BH spin extraction can depend on the disk geometric properties, which in turn are determined by the disk cooling regime, as we now discuss.
@@
BH accretion rate $\dot{m}$ and how close it is to the Eddington value $\dot{m}_{\mathrm{Edd}}$, at which the radiation becomes as important as gravity, roughly sets the disk cooling regime [e.g., 102]. In the highly super-Eddington regime ($\dot{m}\gg\dot{m}_{\mathrm{Edd}}$), the disk is radiatively inefficient: photons are trapped and advected with the gas [74, 100, 82, 31]. This results in a hot accretion flow, which takes the form of a geometrically thick accretion disk. In numerical simulations, such disks typically have a thermal scale height of $h/r\simeq 0.3-0.5$ [6]. The extremely sub-Eddington regime, $\dot{m}\ll 0.01\dot{m}_{\mathrm{Edd}}$, also leads to a thick radiatively inefficient accretion flow (RIAF), which takes the form of a thick disk [67, 68, 69]. In the middle, $0.01\dot{m}_{\mathrm{Edd}}\lesssim\dot{m}\lesssim\dot{m}_{\mathrm{Edd}}$, lies the moderately sub-Eddington regime with an accretion disk that radiates efficiently and cools to scale height values that can become very small, $h/r\ll 1$ [86, 70].
@@
In XRBs, the Eddington ratio of the BH correlates with the accretion spectral state [16, 17, 76, 26]. The soft state, with a predominantly thermal spectrum, is associated with razor-thin disks, whose inner edge is at (or close to) the innermost stable circular orbit [ISCO, 75, 86, 62, 57]. The low/hard state, with a nonthermal power-law spectrum, is linked to thicker disks in the extremely sub-Eddington regime, with electrons that are much cooler than the ions [58, 57, 51].
@@
AGN also come in a wide range of Eddington ratios. Generally, we expect luminous disks to be geometrically thinner than radiatively inefficient disks, although the iron opacity bump and magnetic support may complicate the picture [e.g., 37, 38, 29, 88, 39]. About $10\%$ of luminous quasars are radio-loud, which seems to be related to the strength of their jets [42, 33, 41]. Like XRBs, the measured AGN BH spins are often large [80].
@@
More extreme accretion systems, such as collapsing massive stars producing BHs (collapsars) and binary compact object mergers, can achieve ultra-high super-Eddington accretion rates at which neutrino cooling can reduce their aspect ratio to $h/r\sim 0.2$ [7].
@@
This variety of disk thickness values across different types of astrophysical sources motivates a systematic exploration of the effects that the disk thickness has on the BH spin evolution. For this, it is essential to account for gas accretion and ejection physics, including the effects of strong magnetic field, radiation, and curved spacetime. This ambitious goal can be achieved with general relativistic (radiation) magnetohydrodynamic (GR(R)MHD) simulations. In this work, we focus on magnetically arrested disks (MADs), which are the natural end state of a disk with a sufficiently large vertical magnetic flux reservoir [95, 91, 34]. Such highly magnetized disk states have also been found to emerge naturally due to dynamo action within the accretion disk [52, 36]. MADs can also form in transient sources whose mass accretion rate decreases over time, e.g., TDEs [93, 43] and compact binary mergers [18, 9, 28].
@@
BH equilibrium spin is defined as the BH spin for which the angular momentum lost through the jet balances the angular momentum and mass supplied by the disk. For thick, nonradiative MADs, GRMHD simulations revealed that the equilibrium spin is small, $a_{\mathrm{eq,MAD}}^{\mathrm{thick}}\simeq 0.035\text{--}0.07$ [92, 96, 65, 54]. Lowell et al. [54] found that the BH only needs to accrete $20\%$ of its initial mass to spin down from $a=1$ to $a=0.2$. They also found that the BH loses most of its spin energy to the jets and that the positive angular momentum contribution from the accretion flow is suppressed by the large-scale electromagnetic torques in the accretion disk (see also [56]). They then used this insight to derive a semi-analytic model for spin evolution in the thick nonradiative MAD regime. Using this model, they predicted that luminous, thin MADs would have low equilibrium spin values, $a_{\mathrm{eq}}\sim 0.3\text{--}0.5$. This is much lower than what is obtained by Ricarte et al. [81] based on their GRMHD simulations ($a_{\mathrm{eq}}\sim 0.8$), expected in NT disks ($a_{\mathrm{eq}}\sim 1$), and measured in weakly magnetized “standard and normal evolution” (SANE) accretion disks ($a_{\mathrm{eq}}\sim 0.9$) [25, 6].
@@
In this work, we perform high-resolution 3D GRMHD simulations of cooled BH accretion disks in the MAD state to study BH spin evolution for thin accretion disks. This allows us to extend the work of Lowell et al. [54] from the radiatively inefficient thick to luminous MADs of varying disk scale height to physically interpret our simulation results.
@@
Section II.1 presents the BH spin evolution equations. Section II.2 discusses the standard Novikov-Thorne disk model [73]. Section II.3 describes our numerical setup and simulations. Section II.4 explains how we calculate the torques on the BH. Section III lays out our key results on spin evolution in thin MADs. Section IV derives our new thin MAD spin evolution model and spin-down timescales. Section V explores the dynamic link between the jet and disk magnetic structures, and their effect on the spin-down. Finally, Section VI discusses the results and concludes. Throughout, we use units such that $G=c=1$. We adopt Lorentz-Heaviside units for the magnetic field, so that we absorb the factor of $(4\pi)^{-1/2}$ into the definition of the magnetic field. As a result, the magnetic pressure is $p_{\mathrm{mag}}=b^{2}/2$, where $b$ is the comoving magnetic field strength.
II Methods and Equations
II.1 Spin evolution equations
We define the dimensionless BH spin as
@@ II.1 Spin evolution equations
where $J$ is the angular momentum and $M$ is the mass of the BH, so that it spans $1\leq a\leq 1$. We can then compute the spin evolution of the BH by solving a set of coupled ODEs for the total energy,
$dM$ $=e_{\mathrm{in}}dm$ (2)
and angular momentum,
$dJ$ $=l_{\mathrm{in}}dm,$ (3)
accreted by the BH [63, 54], where $l_{\mathrm{in}}$ and $e_{\mathrm{in}}$ are the specific angular momentum and energy fluxes, respectively, on the BH horizon, $r_{\mathrm{H}}=r_{\mathrm{g}}(1+\sqrt{1-a^{2}})$, and $dm$ is the amount of accreted mass. Here, $r_{\mathrm{g}}=GM/c^{2}$ is the BH gravitational radius. Combining Equations (2) and (3) gives us the ODE for the change in BH spin with BH mass,
$\frac{da}{d\ln M}=\frac{l_{\mathrm{in}}}{e_{\mathrm{in}}}-2a=\frac{s}{e_{\mathrm{in}}},$ (4)
where $\ln$ is the natural logarithm, and we have defined the dimensionless spin-up parameter,
$s\equiv\frac{da}{dm}M=l_{\mathrm{in}}-2ae_{\mathrm{in}},$ (5)
which roughly tells us the change in spin after the BH accretes its own mass (i.e., $dm\sim M$). Computing the BH spin evolution is a question of determining, as a function of BH spin, the specific fluxes, $l_{\mathrm{in}}$ and $e_{\mathrm{in}}$, both of which depend on the (to be determined) accretion flow and jet structures.
II.2 Novikov-Thorne disk model
The NT disk model [73] is an analytic model for a razor-thin accretion disk where the orbits are Keplerian, and the accretion mechanism is a viscosity of turbulent origin. The specific energy and angular momentum fluxes measured at the BH event horizon are [2, 63],
$e_{\mathrm{in}}$ $=\left(1-\frac{2r_{\mathrm{g}}}{3R_{\mathrm{ISCO}}}\right)^{1/2}$ (6)
@@ II.2 Novikov-Thorne disk model
$l_{\mathrm{in}}$ $=\frac{2M}{3^{3/2}}\left[1+2(3R_{\mathrm{ISCO}}/r_{\mathrm{g}}-2)^{1/2}\right],$ (7)
where $R_{\mathrm{ISCO}}$ is the radius of the innermost stable circular orbit [ISCO, 87],
$R_{\mathrm{ISCO}}/r_{\mathrm{g}}$ $=3+Z_{2}-\operatorname{sign}(a)\left[(3-Z_{1})(3+Z_{1}+2Z_{2})\right]^{1/2},$ (8)